# Depth and slip ratio dependencies of friction for a sphere rolling on a granular slope

### v1

@@ 5 Discussion
Fig. 9 shows that $\mu_{\mathrm{d}}$ generally decreases with increasing $\alpha$. This tendency might be qualitatively explained as follows. Figure 8 shows that the size of bump is larger during rolling down motion than during rolling up motion, resulting in a larger contact area i.e., larger $\mu_{\mathrm{d}}$. The relationship between contact area and resistance is qualitatively consistent with the Bekker–Wong–Reece model [30] in terramechanics. In terramechanics models [30], the resistance force exerted on a wheel with radius $R$ and width $b$ is given by $Rb\int_{\theta_{\mathrm{r}}}^{\theta_{\mathrm{f}}}\sigma(\theta)\sin\theta\,d\theta$, where $\theta_{\mathrm{f}}$ and $\theta_{\mathrm{r}}$ are the entry and exit angles, respectively, and $\sigma(\theta)$ is the normal stress, as shown in Fig. 10. The relationship between $\sigma(\theta)$ and the sinking depth $h(\theta)$ is empirically known to follow $\sigma(\theta)\propto h(\theta)^{n}$ with $n>0$. This empirical relation indicates that, as $\theta_{\mathrm{f}}$ increases due to a larger contact area, the resistance force also increases. Accordingly, in the present study, $\mu_{\mathrm{d}}$ is expected to increase with increasing contact area due to bump formation at negative slope angle $\alpha$. While the previous study [8] focused on the decelerating motion of a rolling sphere over a narrow range of slope angles, this study reveals the dependence of $\mu_{\mathrm{d}}$ on $\alpha$ by examining a wide range of $\alpha$ values in both the uphill [10] and downhill directions.
@@
As shown in Fig. 11, $\mu_{\mathrm{d}}$ roughly decreases linearly with $s$ as an overall trend. This result suggests that, in the regime of high slip ratio $s$, rotational motion is partly converted into a driving force for translational motion. Seguin conducted numerical simulations of the drag force exerted on a wedge-shaped intruder moving in a granular medium [26]. Their results showed that the drag force decreases and rotational torque increases with increasing rotational speed, indicating a transfer of mechanical power from translational motion to rotational motion. A similar mechanism may also be present in the present experiment.
As shown in Fig. 4, $\delta$ increases with $\rho_{\mathrm{s}}$. In addition, Fig. 9 shows that $\mu_{\mathrm{d}}$ also increases roughly with $\rho_{\mathrm{s}}$. Therefore, we investigate the relationship between $\mu_{\mathrm{d}}$ and $\delta/R$. It is difficult to obtain a clear relation between $\mu_{\mathrm{d}}$ and $\delta/R$ for each individual angle $\alpha$ because of the limitations in measurement accuracy. To address this, we divide $\alpha$ into three groups that exhibit similar values of $\mu_{\mathrm{d}}$: $\alpha=-20^{\circ},-15^{\circ},-10^{\circ}$], $\alpha=-5^{\circ},0^{\circ},5^{\circ}$], and $\alpha=10^{\circ},15^{\circ},20^{\circ}$]. For each group, we compute the average values of $\mu_{\mathrm{d}}$ and $\delta/R$ and analyze the relationship between them, as shown in Fig. 12.
@@
Specifically, we investigate the relationship between $\mu_{0},(=\mu_{\mathrm{d}}-\beta(\delta/R))$ and $s$, as shown in Fig. 13. The value of $\mu_{0}$ decreases linearly with $s$, as indicated in Fig. 13. By applying least-square fitting to the data in Fig. 13, we obtain the empirical relation, $\mu_{\mathrm{d}}-\beta(\delta/R)=0.32-0.44s~(\beta=0.41)$. To the best of our knowledge, this study provides the first quantitative empirical formula linking $\mu_{\mathrm{d}}$, $s$, and $\delta/R$. This relation suggests that $\mu_{\mathrm{d}}$ consists of two contributions: one associated with the sinking deformation and the other associated with the bump formation in front of the sphere, which is related to the slip ratio $s$. The frictional contribution due to sinking is determined solely by the density of the sphere and is independent of the angle $\alpha$. In contrast, the contribution associated with bump formation correlates with the slip ratio $s$, which depends on the angle $\alpha$.
@@
Figure 8: The image of the sphere (a) rolling down at $\alpha\simeq-20^{\circ}$, and (b) rolling up at $\alpha\simeq 20^{\circ}$ from Fukumoto et al [10]. The radius of polyethylene sphere is 6.35 mm in both (a) and (b). The yellow curve indicates the edge of the bump.
Figure 9: The relation between $\mu_{\mathrm{d}}$ and angle $\alpha$. $\alpha<0$ and $\alpha>0$ correspond to rolling down and up cases respectively. We use the data ($\alpha>0$) obtained by Fukumoto et al. [10]. Error bars indicate the standard deviation of all the data.
@@
Figure 11: The relation between $\mu_{\mathrm{d}}$ and slip ratio $s$, except for the glass sphere, at each $\alpha$. Error bars indicate the standard deviation of all the data. The data of rolling up is taken from Fukumoto et al. [10].
Figure 12: The relation between $\mu_{\mathrm{d}}$ and $\delta/R$. Error bars indicate the standard deviation of all the data. Filled blue, empty blue, and filled red symbols represent the data rolling down the granular slop for $\alpha=-20^{\circ},-15^{\circ},-10^{\circ}$], $\alpha=-5^{\circ},0^{\circ},5^{\circ}$], $\alpha=10^{\circ},15^{\circ},20^{\circ}$], respectively. Blue solid and dashed lines indicate the least-square linear fitting line [Eq. (6)] with the slope of $\beta=0.41\pm 0.04$ and $0.45\pm 0.04$ with the intercept of $\mu_{0}=0.32\pm 0.02$ and $0.1\pm 0.02$ respectively. The red dashed line indicates the least-square fitting line with the slope of $\beta=0.36\pm 0.1$ with the intercept of $\mu_{0}=0.03\pm 0.07$. The fitting error is the standard deviation of the least-square fitting. We use the data ($\alpha>0$) obtained from [10].
Figure 13: The relation between $\mu_{\mathrm{d}}-\beta(\delta/R)$ and $s$ at each $\alpha$ ($\beta=0.41$). The black solid line indicates the least-square linear fitting line, $\mu_{\mathrm{d}}-\beta(\delta/R)=0.32-0.44s$. Error bars indicate the standard deviation of all the data.
@@ 6 Conclusion
In this study, we characterize the friction of a sphere rolling down a granular slope. The dynamics are experimentally investigated by varying the initial velocity, slope angle, and sphere density. The sinking depth scales as, $\delta/R=C_{\rho}\,(\rho_{\mathrm{s}}/\rho_{\mathrm{g}})^{3/4},$ which is consistent with previous results for uphill rolling [10]. The translational motion exhibits constant deceleration. To characterize the motion, we introduce the effective friction coefficient $\mu_{\mathrm{d}}$. $\mu_{\mathrm{d}}$ shows a decreasing tendency with the slope angle $\alpha$ and the slip ratio $s$. Also, $\mu_{\mathrm{d}}$ increases linearly $\delta/R$, $\mu_{\mathrm{d}}=\beta(\delta/R)+\mu_{0}$. $\beta\simeq 0.41$ does not vary significantly, while the value of $\mu_{0}$ is linearly decreasing with $s$. Understanding why $\mu_{\mathrm{d}}$ takes larger values in the downhill direction, while maintaining almost the same proportional slope $\beta$, remains a crucial issue for future work.
[…]
@@
Sample text inserted for demonstration.
@@
Sample text inserted for demonstration.
[…]
Sample text inserted for demonstration.

### v2

@@ 5 Discussion
Fig. 9 shows that $\mu_{\mathrm{d}}$ generally decreases with increasing $\alpha$ as an overall trend. From the correlation degree characterized by the correlation coefficient, $\mu_{\mathrm{d}}$ has a clear negative correlation with $\alpha$. This tendency might be qualitatively explained as follows. Figure 8 shows that the size of bump is larger during rolling down motion than during rolling up motion, resulting in a larger contact area i.e., larger $\mu_{\mathrm{d}}$. The relationship between contact area and resistance is qualitatively consistent with the Bekker–Wong–Reece model [31] in terramechanics. In terramechanics models [31], the resistance force exerted on a wheel with radius $R$ and width $b$ is given by $Rb\int_{\theta_{\mathrm{r}}}^{\theta_{\mathrm{f}}}\sigma(\theta)\sin\theta\,d\theta$, where $\theta_{\mathrm{f}}$ and $\theta_{\mathrm{r}}$ are the entry and exit angles, respectively, and $\sigma(\theta)$ is the normal stress, as shown in Fig. 10. The relationship between $\sigma(\theta)$ and the sinking depth $h(\theta)$ is empirically known to follow $\sigma(\theta)\propto h(\theta)^{n}$ with $n>0$. This empirical relation indicates that, as $\theta_{\mathrm{f}}$ increases due to a larger contact area, the resistance force also increases. Accordingly, in the present study, $\mu_{\mathrm{d}}$ is expected to increase with increasing contact area due to bump formation at negative slope angle $\alpha$. While the previous study [8] focused on the decelerating motion of a rolling sphere over a narrow range of slope angles, this study reveals the dependence of $\mu_{\mathrm{d}}$ on $\alpha$ by examining a wide range of $\alpha$ values in both the uphill [10] and downhill directions.
@@
As shown in Fig. 11, $\mu_{\mathrm{d}}$ roughly decreases linearly with $s$ as an overall trend. The correlation coefficient values suggest the clear negative correlation between $\mu_{\mathrm{d}}$ and $s$. This result suggests that, in the regime of high slip ratio $s$, rotational motion is partly converted into a driving force for translational motion. Seguin conducted numerical simulations of the drag force exerted on a wedge-shaped intruder moving in a granular medium [27]. Their results showed that the drag force decreases and rotational torque increases with increasing rotational speed, indicating a transfer of mechanical power from translational motion to rotational motion. A similar mechanism may also be present in the present experiment.
As shown in Fig. 4, $\delta$ increases with $\rho_{\mathrm{s}}$. In addition, Fig. 9 shows that $\mu_{\mathrm{d}}$ also increases roughly with $\rho_{\mathrm{s}}$. Therefore, we investigate the relationship between $\mu_{\mathrm{d}}$ and $\delta/R$. It is difficult to obtain a clear relation between $\mu_{\mathrm{d}}$ and $\delta/R$ for each individual angle $\alpha$ because of the limitations in measurement accuracy. To address this, we divide $\alpha$ into three groups that exhibit similar values of $\mu_{\mathrm{d}}$: $\alpha=-20^{\circ},-15^{\circ},-10^{\circ}$], $\alpha=-5^{\circ},0^{\circ},5^{\circ}$], and $\alpha=10^{\circ},15^{\circ},20^{\circ}$]. For each group, we compute the average values of $\mu_{\mathrm{d}}$ and $\delta/R$ and analyze the relationship between them, as shown in Fig. 12. The correlation coefficient values show that $\mu_{\mathrm{d}}$ has a clear positive correlation with $\delta/R$.
@@
Specifically, we investigate the relationship between $\mu_{0},(=\mu_{\mathrm{d}}-\beta(\delta/R))$ and $s$, as shown in Fig. 13. The correlation coefficient values show that $\mu_{0}$ has a clear negative correlation with $s$. The value of $\mu_{0}$ decreases linearly with $s$, as indicated in Fig. 13. By applying least-square fitting to the data in Fig. 13, we obtain the empirical relation, $\mu_{\mathrm{d}}-\beta(\delta/R)=0.32-0.44s~(\beta=0.41)$. To the best of our knowledge, this study provides the first quantitative empirical formula linking $\mu_{\mathrm{d}}$, $s$, and $\delta/R$. This relation suggests that $\mu_{\mathrm{d}}$ consists of two contributions: one associated with the sinking deformation and the other associated with the bump formation in front of the sphere, which is related to the slip ratio $s$. The frictional contribution due to sinking is determined solely by the density of the sphere and is independent of the angle $\alpha$. In contrast, the contribution associated with bump formation correlates with the slip ratio $s$, which depends on the angle $\alpha$.
@@
In this study, our objective is to investigate the interaction between a rolling object and deformable granular media by simplifying the system and eliminating complex phenomena. To achieve this, we use dry grains in order to exclude the effects of adhesion and grain aggregation. In addition, we employed spheres with nearly identical surface conditions (i.e., similar static friction coefficients), so that sphere density would serve as the primary variable rather than surface properties. We focus on macroscopic interactions as a first step toward understanding the mechanics between a sphere and granular media. A more detailed investigation at the microscopic level would further advance understanding of the underlying interaction mechanisms. Although the present study intentionally eliminates complex factors, it would be interesting in future work to examine the effect of surface conditions by using actively adhering spheres, or to investigate how changes in substrate properties such as particle density and size distribution affect the dynamics.
In this research, it is assumed that the rigidity of the spheres and granular materials does not affect the behavior of the spheres, but the case where the rigidity of each material has an effect is also an interesting issue that should be solved in the future.
@@
Figure 8: The image of the sphere (a) rolling down at $\alpha\simeq-20^{\circ}$, and (b) rolling up at $\alpha\simeq 20^{\circ}$ from Fukumoto et al [10]. The radius of polyethylene sphere is 6.35 mm in both (a) and (b). The yellow curve indicates the edge of the bump. The slip ratio $s$ is smaller in (a) and larger in (b).
Figure 9: The relation between $\mu_{\mathrm{d}}$ and angle $\alpha$. $\alpha<0$ and $\alpha>0$ correspond to rolling down and up cases respectively. We use the data ($\alpha>0$) obtained by Fukumoto et al. [10]. Error bars indicate the standard deviation of all the data. R=-0.97(polyethylene), -0.95(polyacetal), -0.92(glass), -0.87(alumina ceramic), R: the correlation coefficient.
@@
Figure 11: The relation between $\mu_{\mathrm{d}}$ and slip ratio $s$, except for the glass sphere, at each $\alpha$. Error bars indicate the standard deviation of all the data. The data of rolling up is taken from Fukumoto et al. [10]. R=-0.96(polyethylene), -0.94(polyacetal), -0.82(alumina ceramic), R: the correlation coefficient.
Figure 12: The relation between $\mu_{\mathrm{d}}$ and $\delta/R$. Error bars indicate the standard deviation of all the data. Filled blue, empty blue, and filled red symbols represent the data rolling down the granular slop for $\alpha=-20^{\circ},-15^{\circ},-10^{\circ}$], $\alpha=-5^{\circ},0^{\circ},5^{\circ}$], $\alpha=10^{\circ},15^{\circ},20^{\circ}$], respectively. Blue solid and dashed lines indicate the least-square linear fitting line [Eq. (6)] with the slope of $\beta=0.41\pm 0.04$ and $0.45\pm 0.04$ with the intercept of $\mu_{0}=0.32\pm 0.02$ and $0.1\pm 0.02$ respectively. The red dashed line indicates the least-square fitting line with the slope of $\beta=0.36\pm 0.1$ with the intercept of $\mu_{0}=0.03\pm 0.07$. The fitting error is the standard deviation of the least-square fitting. We use the data ($\alpha>0$) obtained from [10]. R=0.75 $\alpha=-20^{\circ},-15^{\circ},-10^{\circ}$], 0.83[$\alpha=-5^{\circ},0^{\circ},5^{\circ}$], 0.84[$\alpha=10^{\circ},15^{\circ},20^{\circ}$], R: the correlation coefficient.
Figure 13: The relation between $\mu_{\mathrm{d}}-\beta(\delta/R)$ and $s$ at each $\alpha$ ($\beta=0.41$). The black solid line indicates the least-square linear fitting line, $\mu_{\mathrm{d}}-\beta(\delta/R)=0.32-0.44s$. Error bars indicate the standard deviation of all the data. The correlation coefficient R, R=-0.78.
@@ 6 Conclusion
In previous studies, no universal law was established for the relationship between effective friction and sinking depth that applies to both rolling-up and rolling-down cases. In this study, we characterize the friction of a sphere rolling down a granular slope. The dynamics are experimentally investigated by varying the initial velocity, slope angle, and sphere density. We selected a parameter range that was not explored in previous studies of rolling-down motion. The sinking depth scales as, $\delta/R=C_{\rho}\,(\rho_{\mathrm{s}}/\rho_{\mathrm{g}})^{3/4},$ which is consistent with previous results for uphill rolling [10]. The translational motion exhibits constant deceleration. To characterize the motion, we introduce the effective friction coefficient $\mu_{\mathrm{d}}$. $\mu_{\mathrm{d}}$ shows a decreasing tendency with the slope angle $\alpha$ and the slip ratio $s$. Also, $\mu_{\mathrm{d}}$ increases linearly $\delta/R$, $\mu_{\mathrm{d}}=\beta(\delta/R)+\mu_{0}$. $\beta\simeq 0.41$ does not vary significantly, while the value of $\mu_{0}$ is linearly decreasing with $s$. Understanding why $\mu_{\mathrm{d}}$ takes larger values in the downhill direction, while maintaining almost the same proportional slope $\beta$, remains a crucial issue for future work.
@@
Further information is available in the online guidelines: https://doi.org/10.5281/zenodo.18810364
@@
The typical examples of actual movies